# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
Since $p^{-1}(a)$ is invariant under permuting $x_{i}$-factors, we set $i=1$ for simplicity. Take a general $a\in\sigma_{k}(Z)={\mathbb{P}}^{\beta}$. Since $\operatorname{Vertex}(\sigma_{k-1}(Z))$ is a linear subvariety of $\sigma_{k-1}(Z)\subsetneq{\mathbb{P}}^{\beta}$ and $Z$ is non-degenerate in ${\mathbb{P}}^{\beta}$, $\operatorname{Vertex}(\sigma_{k-1}(Z))\cap Z\subsetneq Z$. This implies that for general $b\in Z$, $\dim\langle\,b,\sigma_{k-1}(Z)\,\rangle>\dim(\sigma_{k-1}(Z))$ so that $\langle\,b,\sigma_{k-1}(Z)\,\rangle={\mathbb{P}}^{\beta}$. Thus, any given general point $a\in{\mathbb{P}}^{\beta}$ sits on a line $\langle\,b,c\,\rangle$ for any general $b\in Z$ and for some general $c\in\sigma_{k-1}(Z)$ and $(a,b,\tilde{x_{2}},\dots,\tilde{x_{k}})\in p^{-1}(a)$, which means $q_{1}(p^{-1}(a))=Z$. ∎
Now let us focus on the case $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta=\binom{m+d}{m}-1}$, the image of the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$. Here we prove a very useful proposition, which has an independent interest itself. In Proposition 15, we consider the entry locus of a general point in $Z$ and estimate the dimension of the linear span of its image under ($d-1$)-uple Veronese embedding $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}=\binom{m+d-1}{m}-1}$.
@@
Let $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta}$ with $d\geq 3$, $22\leq m\leq k-$ 2, and $\beta=\binom{m+d}{m}-1<km+k-1$. Assume $\dim(\sigma_{k-1}(Z))=(k-1)m+k-2<\beta$ (in other words, the ($k-1$)-th secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$). Let $J\subset{\mathbb{P}}^{\beta}\times(Z)^{k}$ be the Zariski closure of the incidence (5), let $(a,x_{1},\dots,x_{k})\in J$ be a general point, and let $F\subset J$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$. Then the following holds.
@@
    If $(k-1)m+k=\binom{m+d}{m}$, then it holds $q_{i}(F)=Z$. In addition, if $(d,m)\neq(3,2)$, then $\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle\geq k+m$.
@@
    Let $\mu=\frac{\binom{m+d}{m}}{m+1}$. Then inequalities $\beta<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ (i.e., $\binom{m+d}{m}<km+k$ and $(k-1)m+k-1<\binom{m+d}{m}$) are equivalent to $\mu<k<\mu+1$; it happens if and only if $\mu\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$.
@@
Let $d,m,k$ be integers such that $d\geq 3$, $22\leq m\leq k-$ 2, and $\binom{m+d}{m}-1<km+k-1$.
@@
    If $(k-1)m+k<\binom{m+d}{m}$, then $\binom{m+d-1}{m}-1-2m-k\geq 0$.
@@
Proof.
(i) From Remark 18 (a) below, we may assume $m\geq 3$. Let $\delta:=\binom{m+d-1}{m}-(1+k+2m)$. Since $\binom{m+d-1}{m}=\frac{d}{m+d}\binom{m+d}{m}$, using $(k-1)m+k+1\leq\binom{m+d}{m}$, we have
$\delta=\binom{m+d-1}{m}-(1+k+2m)\geq\frac{1}{m+d}\Bigl{(}d((k-1)m+k+1)-(m+d)(1+k+2m)\Bigr{)},$
Setting $m=k-a$ with $a\geq 2$, we have
(7) $d((k-1)m+k+1)-(m+d)(1+k+2m)=km(d-1)-3dm-m-2m^{2}\\ =(m+a)m(d-1)-3dm-m-2m^{2}=m(dm+(a-3)d-3m-(a+1))\\ \geq m((d-3)(m-1)-6)$
Then, $\delta\geq 0$ holds in the following three cases: $d\geq 6$ and $m\geq 3$; $d=5$ and $m\geq 4$; or $d=4$ and $m\geq 7$. In addition, in Remark 18 (b) (see also Remark 16 (a)), we explicitly check $\delta\geq 0$ if $d=4$ and $m\leq 6$, and check that there is no $k$ in our range if $d=5$ and $m=3$.
On the other hands, when $d=3$, (7) implies
$d((k-1)m+k+1)-(m+d)(1+k+2m)=m((a-3)3-(a+1))=m(2a-10).$
Hence, $\delta\geq 0$ holds if $d=3$ and $a\geq 5$. For $d=3$ and $m=k-a$ with $a=2,3,4$, we have
@@
(ii) Let $\delta:=\binom{m+d-1}{m}-(1+k+m)$. In the same way as (i), using $(k-1)m+k=\binom{m+d}{m}$, we have
@@
On the other hands, if $d=3$ and $m=3$, then $(k-1)m+k=\binom{m+d}{m}$ implies $204k-3=\frac{(d+3)(d+2)(d+1)}{3!}=$ 20; this case does not occur since $k$ cannot be an integer.
@@
Remark 18.
In the proof of Lemma 17, we refer to two calculations as follows.
  1. (a)
    Let $m=2$ and $km+k-1>\binom{m+d}{m}-1$. Then $(k-1)m+k<\binom{m+d}{m}$ does not occur. Otherwise, we have $3k-2<\binom{m+d}{m}=\frac{(d+2)(d+1)}{2}<3k$, and then $\frac{(d+2)(d+1)}{2}=3k-1$. Considering the congruence modulo $3$, we have $(d+2)(d+1)\equiv 6k-2\equiv 1\pmod{3}$. Then the possible values of $(d+2)(d+1)$ are
    modulo $3$, which is absurd.
  2. (b)
    For $d=3,4,5$, we calculate numbers $m$ such that $\mu:=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$ satisfies $(k-1)m+k<\beta$. For $d=3$, the smallest $m=5$. For $d=4$, the smallest $m=3$ and the next smallest $m=7$. For $d=5$, the smallest $m=9$. The explicit values of $\delta=\binom{m+d-1}{m}-(1+k+2m)$ for them are obtained as follows:
@@
Let $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ be general $k$ points, let $v_{e}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{e}=\binom{m+e}{m}-1}$ be the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$, and let $R\subset{\mathbb{P}}^{\beta_{e}}$ be any $k$-plane containing the ($k-1$)-plane $M:=\langle\,v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$. Then the following holds.
@@
    $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0~{},$
    where the left-hand side means dimension of component(s) passing through $v_{e}(x_{1}^{\prime})$. In particular, the set of a point $\left\{v_{e}(x_{1}^{\prime})\right\}$ is an irreducible component of ${R\cap v_{e}({\mathbb{P}}^{m})}$.
@@
be the linear projection from the ($k-2$)-plane $\langle\,v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$. If $k\leq\beta_{e}-m$, then the generalized trisecant lemma ([28, proposition 1.4.3]) implies $M\cap v_{e}({\mathbb{P}}^{m})=\left\{v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\right\}$; in particular, $C\not\subset M$. Hence $C$ is not contracted to a point under $\pi$. Then $\overline{\pi(C)}\subset{\mathbb{P}}^{\beta_{e}-k+1}$ is a curve passing through $\pi(v_{e}(x_{1}^{\prime}))$, and then $\overline{\pi(C)}=\overline{\pi(R)}$, which is indeed a line. It follows
$\overline{\pi(C)}=\overline{\pi(R)}\subset\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(v_{e}({\mathbb{P}}^{m}))}$
@@
from the $m$-plane ${\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}\subset{\mathbb{P}}^{\beta_{e}}$, and its restriction $\tilde{\pi}=\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}|_{v_{e}({\mathbb{P}}^{m})}$ on the Veronese variety. Note that $v_{e}({\mathbb{P}}^{m})\cap{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}=\left\{v_{e}(x_{1}^{\prime})\right\}$. If $R\subset{\mathbb{P}}^{\beta}$ satisfies (8), we have
@@
Let $e\geq 3$ and $k\leq\beta_{e}-2m$. Suppose that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})>0$, which means the existence of a curve $C$ satisfying the condition of (i). Since $\operatorname{codim}(\tilde{\pi}(v_{e}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{e}-m-1})-(k-1)\geq(\beta_{e}-2m-1)-(k-1)\geq 0$, again the trisecant lemma implies that the right hand side in (9) is only the set of $k-1$ points ${\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))}$. Thus, each irreducible subset $D\subset R\cap v_{e}({\mathbb{P}}^{m})$ satisfies
@@
for some $l=2,\dots,k$. (At least, taking $D=C$, we exactly have (10).) From $e\geq 3$, $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))=2m+1$ (i.e., non-defective). Then $\tilde{\pi}$ must be a generically finite map since general two tangent spaces of $\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$ do not meet each other, and we reach a contradiction since $v_{e}(x_{l}^{\prime})$ is a general point. It implies that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0$.
Finally, we consider the case of $e=2$. Then it is well known that the tangential projection $\tilde{\pi}:v_{2}({\mathbb{P}}^{m})\dashrightarrow\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))$ has fibers of dimension $1$. Moreover, as in Remark 8 (a) we know that $\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{2}=\frac{(m+2)(m+1)}{2}-1}\dashrightarrow{\mathbb{P}}^{\beta_{2}-m-1=\frac{(m+1)m}{2}-1}$ satisfies the commutative diagram:

### v3

@@ 2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
Since $p^{-1}(a)$ is invariant under permuting $x_{i}$-factors, we set $i=1$ for simplicity. Take a general $a\in\sigma_{k}(Z)={\mathbb{P}}^{\beta}$. Since $\operatorname{Vertex}(\sigma_{k-1}(Z))$ is a linear subvariety of $\sigma_{k-1}(Z)\subsetneq{\mathbb{P}}^{\beta}$ and $Z$ is non-degenerate in ${\mathbb{P}}^{\beta}$,
$\operatorname{Vertex}(\sigma_{k-1}(Z))\cap Z\subsetneq Z.$
This implies that for general $x\in Z$, $\dim\langle\,x,\sigma_{k-1}(Z)\,\rangle>\dim(\sigma_{k-1}(Z))$ so that $\langle\,x,\sigma_{k-1}(Z)\,\rangle={\mathbb{P}}^{\beta}$. Thus, any given general point $a\in{\mathbb{P}}^{\beta}$ sits on a line $\langle\,x,b\,\rangle$ for any general $x\in Z$ and for some general $b\in\sigma_{k-1}(Z)$. Taking $k-1$ points $\tilde{x_{2}},\dots,\tilde{x_{k}}\in Z$ such that $b\in\langle\,\tilde{x_{2}},\dots,\tilde{x_{k}}\,\rangle$, we have $(a,x,\tilde{x_{2}},\dots,\tilde{x_{k}})\in p^{-1}(a)$, which means $q_{1}(p^{-1}(a))=Z$. ∎
Now let us focus on the case $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta=\binom{m+d}{m}-1}$, the image of the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$. Here we prove a very useful proposition, which is of independent interest itself. In Proposition 15, we consider the entry locus of a general point in $Z$ and estimate the dimension of the linear span of its image under ($d-1$)-uple Veronese embedding $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}=\binom{m+d-1}{m}-1}$.
@@
Let $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta}$ with $d\geq 3$, $22\leq m\leq k-$ 2, and $\beta=\binom{m+d}{m}-1<km+k-1$. Assume $\dim(\sigma_{k-1}(Z))=(k-1)m+k-2<\beta$ (in other words, the ($k-1$)-th secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$), where we have $(k-1)m+k\leq\binom{m+d}{m}$. Let $J\subset{\mathbb{P}}^{\beta}\times(Z)^{k}$ be the Zariski closure of the incidence (9), let $(a,x_{1},\dots,x_{k})\in J$ be a general point, and let $F\subset J$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$. Then the following holds.
@@
    If $(k-1)m+k=\binom{m+d}{m}$, then $q_{i}(F)=Z$. In addition, if $(d,m)\neq(3,2)$, then $\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle\geq k+m$.
@@
    Two inequalities $\beta=\binom{m+d}{m}-1<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ are equivalent to $\frac{\binom{m+d}{m}}{m+1}<k<\frac{\binom{m+d}{m}}{m+1}+1$; it occurs if and only if $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$.
@@
Let $d,m,k$ be integers such that $d\geq 3$, $22\leq m\leq k-$ 2.
@@
    If $(k-1)m+k<\binom{m+d}{m}<km+k$, then $\binom{m+d-1}{m}-1-2m-k\geq 0$.
@@
  3. (iii)
    If $\binom{m+d}{m}=km+k$, then $\binom{m+d-1}{m}-1-m-k\geq 0$.
Note that (iii) of Lemma 17 is applied in a discussion of the proof of Part (ii) of Theorem 2, though it is not used in this section.
To show the lemma, we need some calculations as follows.
Remark 18.
(a) Let $m=2$ and $km+k>\binom{m+d}{m}$. Then $(k-1)m+k<\binom{m+d}{m}$ does not occur. Otherwise, we have $3k-2<\binom{m+d}{m}=\frac{(d+2)(d+1)}{2}<3k$, and then $\frac{(d+2)(d+1)}{2}=3k-1$. Considering the congruence modulo $3$, we have $(d+2)(d+1)\equiv 6k-2\equiv 1\pmod{3}$. Then the possible values of $(d+2)(d+1)$ are
modulo $3$, which is absurd.
(b) For $d=3,4,5$, we calculate numbers $m$ satisfying the conditions $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$, $k=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$, and $(k-1)m+k<\binom{m+d}{m}$. For $d=3$, the smallest $m=5$. For $d=4$, the smallest $m=3$ and the next smallest $m=7$. For $d=5$, the smallest $m=9$. The explicit values of $\delta=\binom{m+d-1}{m}-(1+k+2m)$ for them are obtained as follows:
Proof of Lemma 17.
(i) From Remark 18 (a), we may assume $m\geq 3$. Let $\delta=\binom{m+d-1}{m}-(1+k+2m)$. Since $\binom{m+d-1}{m}=\frac{d}{m+d}\binom{m+d}{m}$, using $(k-1)m+k+1\leq\binom{m+d}{m}$, we have
$\delta=\binom{m+d-1}{m}-(1+k+2m)\geq\frac{1}{m+d}\Bigl{(}d((k-1)m+k+1)-(m+d)(1+k+2m)\Bigr{)}.$
Setting $k=m+a$ with $a\geq 2$, we have
(11) $d((k-1)m+k+1)-(m+d)(1+k+2m)=m(k(d-1)-3d-1-2m)\\ =m((m+a)(d-1)-3d-1-2m)\geq m((m+2)(d-1)-3d-1-2m)\\ =m(dm-3m-d-3)=m((d-3)(m-1)-6).$
Then, $\delta\geq 0$ holds in the following three cases: $d\geq 6$ and $m\geq 3$; $d=5$ and $m\geq 4$; or $d=4$ and $m\geq 7$. In addition, in Remark 18 (b) (see also Remark 16 (a)), we explicitly check that $\delta\geq 0$ if $d=4$ and $m\leq 6$, and check that there is no $k$ in our range if $d=5$ and $m=3$.
On the other hand, when $d=3$, (11) implies
$d((k-1)m+k+1)-(m+d)(1+k+2m)=m(2(m+a)-10-2m)=m(2a-10).$
Hence, $\delta\geq 0$ holds if $d=3$ and $a\geq 5$. For $d=3$ and $k=m+a$ with $a=2,3,4$, we have
@@
(ii) Let $\delta=\binom{m+d-1}{m}-(1+k+m)$. In the same way as (i), using $(k-1)m+k=\binom{m+d}{m}$, we have
@@
On the other hand, if $d=3$ and $m=3$, then $(k-1)m+k=\binom{m+d}{m}$ implies
$4k-3=\frac{(d+3)(d+2)(d+1)}{3!}=20;$
this case does not occur since $k$ cannot be an integer.
@@
(iii) Next, we assume $\binom{m+d}{m}=km+k$. Then,
$\delta=\binom{m+d-1}{m}-(1+k+m)=\frac{1}{m+d}\Bigl{(}d(km+k)-(m+d)(1+k+m)\Bigr{)}.$
Using $k=m+a=(m+1)+(a-1)$ with $a\geq 2$, we have
$d(km+k)-(m+d)(1+k+m)\\ =(d-1)km-(m+d)(m+1)=(d-1)m(m+1)+(d-1)m(a-1)-(m+d)(m+1)\\ =(m+1)((d-1)m-(m+d))+(d-1)m(a-1)\\ =(m+1)((d-2)(m-1)-2)+(d-1)m(a-1).$
If $d\geq 3,m\geq 2$ and $(d,m)\neq(3,2)$, then we have $\delta\geq 0$. If $(d,m)=(3,2)$, then since
$(m+1)((d-2)(m-1)-2)+(d-1)m(a-1)=-3+4(a-1)\geq 1,$
we also have $\delta\geq 0$. ∎
@@
Let $k,m\geq 2$. Let $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ be general $k$ points, let $v_{e}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{e}=\binom{m+e}{m}-1}$ be the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$, and let $M=\langle\,v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{e}}$. Then, for any $k$-plane $R\subset{\mathbb{P}}^{\beta_{e}}$ containing the ($k-1$)-plane $M$, the following holds.
@@
    $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0,$
    where the left hand side means dimension of component(s) passing through $v_{e}(x_{1}^{\prime})$. In particular, the set of a point $\left\{v_{e}(x_{1}^{\prime})\right\}$ is an irreducible component of ${R\cap v_{e}({\mathbb{P}}^{m})}$.
@@
be the linear projection from the ($k-2$)-plane $\langle\,v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$. If $k\leq\beta_{e}-m$, then the generalized trisecant lemma ([Rus16, Proposition 1.4.3]) implies
$M\cap v_{e}({\mathbb{P}}^{m})=\left\{v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\right\}.$
In particular, $\dim(M\cap v_{e}({\mathbb{P}}^{m}))=0$ and $C\not\subset M$. We have $\pi(v_{e}(x_{1}^{\prime}))\in\overline{\pi(C)}\subset{\mathbb{P}}^{\beta_{e}-k+1}$ because of $v_{e}(x_{1}^{\prime})\in C$. If $C$ is contracted to a point under $\pi$, then $\overline{\pi(C)}=\pi(v_{e}(x_{1}^{\prime}))$, which means that $C\subset M$, a contradiction. Hence $\overline{\pi(C)}$ must be a curve.
For the $k$-plane $R\subset{\mathbb{P}}^{\beta_{e}}$, which contains the ($k-2$)-dimensional center of $\pi$, the image $\overline{\pi(R)}$ is a line in ${\mathbb{P}}^{\beta_{e}-k+1}$. Thus $\overline{\pi(C)}=\overline{\pi(R)}$. Moreover, it follows
$\overline{\pi(C)}=\overline{\pi(R)}=\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(R)}\subset\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(v_{e}({\mathbb{P}}^{m}))}$
@@
from the $m$-plane ${\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}\subset{\mathbb{P}}^{\beta_{e}}$, and its restriction $\tilde{\pi}=\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}|_{v_{e}({\mathbb{P}}^{m})}$ on $v_{e}({\mathbb{P}}^{m})$. Note that the Veronese variety $v_{e}({\mathbb{P}}^{m})$ and any embedded tangent space to $v_{e}({\mathbb{P}}^{m})$ intersect only at one point; in particular, $v_{e}({\mathbb{P}}^{m})\cap{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}=\left\{v_{e}(x_{1}^{\prime})\right\}$. If $R\subset{\mathbb{P}}^{\beta}$ satisfies (12), we have
@@
By Terracini’s lemma, for general $z\in v_{e}({\mathbb{P}}^{m})$, the linear variety $\langle\,\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m}),\mathbb{T}_{z}v_{e}({\mathbb{P}}^{m})\,\rangle$ coincides with an embedded tangent space to $\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$, and is of dimension $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$. Then,
$\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}(\mathbb{T}_{z}v_{e}({\mathbb{P}}^{m}))=\mathbb{T}_{\tilde{\pi}(z)}\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))\subset{\mathbb{P}}^{\beta_{e}-m-1}$
is of dimension $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))-m-1$. It follows
(14) $\dim\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))=\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))-m-1.$
Let $e\geq 3$ and $k\leq\beta_{e}-2m$. Suppose that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})>0$, which means the existence of a curve $C$ satisfying the condition of (i). Since
$\operatorname{codim}(\tilde{\pi}(v_{e}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{e}-m-1})-(k-1)\geq(\beta_{e}-2m-1)-(k-1)\geq 0,$
again the trisecant lemma implies that the right hand side in (13) is only the set of $k-1$ points ${\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))}$. Thus, each irreducible subset $D\subset R\cap v_{e}({\mathbb{P}}^{m})$ satisfies
@@
for some $l=2,\dots,k$. (At least, taking $D=C$, we exactly have (15).) From $e\geq 3$, $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))=2m+1$ (i.e., non-defective). In this case, by (14), $\dim\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))=m$, i.e., the map $\tilde{\pi}$ must be generically finite. Then we reach a contradiction since $v_{e}(x_{l}^{\prime})$ is a general point. It implies that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0$.
Finally, we consider the case of $e=2$. Then $\dim\sigma_{2}(v_{2}({\mathbb{P}}^{m}))=2m$ for $m\geq 2$ (i.e., defective), and by (14), $\dim\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))=m-1$. It means that the tangential projection $\tilde{\pi}:v_{2}({\mathbb{P}}^{m})\dashrightarrow\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))$ has fibers of dimension $1$. Moreover, as in Remark 8 (a) we know that
$\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{2}=\frac{(m+2)(m+1)}{2}-1}\dashrightarrow{\mathbb{P}}^{\beta_{2}-m-1=\frac{(m+1)m}{2}-1}$
satisfies the commutative diagram: